The management of Novak Company has asked its accounting department to describe the effect upon the company’s financial position and its income statements of accounting for inventories on the LIFO rather than the FIFO basis during 2020 and 2021. The accounting department is to assume that the change to LIFO would have been effective on January 1, 2020, and that the initial LIFO base would have been the inventory value on December 31, 2019. The following are the company’s financial statements and other data for the years 2020 and 2021 when the FIFO method was employed. Financial Position as of 12/31/19 12/31/20 12/31/21 Cash $ 115,200 $166,400 $197,120 Accounts receivable 102,400 128,000 153,600 Inventory 153,600 179,200 225,280 Other assets 204,800 217,600 256,000 Total assets $576,000 $691,200 $832,000 Accounts payable $ 51,200 $ 76,800 $ 102,400 Other liabilities 89,600 102,400 140,800 Common stock 256,000 256,000 256,000 Retained earnings 179,200 256,000 332,800 Total liabilities and equity $576,000 $691,200 $832,000 Income for Years Ended 12/31/20 12/31/21 Sales revenue $1,152,000 $1,728,000 Less: Cost of goods sold 646,400 967,680 Other expenses 262,400 389,120 908,800 1,356,800 Income before income taxes 243,200 371,200 Income taxes (40%) 97,280 148,480 Net income $145,920 $ 222,720 Other data: 1. Inventory on hand at December 31, 2019, consisted of 51,200 units valued at $3.00 each. 2. Sales (all units sold at the same price in a given year): 2020-192,000 units @ $6.00 each 2021-230,400 units @ $7.50 each 3. Purchases (all units purchased at the same price in given year): 2020-192,000 units @ $3.50 each 2021-230,400 units @ $4.40 each 4. Income taxes at the effective rate of 40% are paid on December 31 each year. Name the account(s) presented in the financial statements that would have different amounts for 2021 if LIFO rather than FIFO had been used, and state the new amount for each account that is named.
Answer #1
Workings: FIFO LIFO Diff.
12/31/2019 Ending inventory 51200*3= 153600 153600
2020 Purchases 192000*3.50= 672000 672000
COGAS 825600 825600
2020 Sales/COGS 192000 51200*3= 153600 192000*3.50= 672000
140800*3.5= 492800 25600
12/31/2020 Ending inventory 51200*3.50 179200 51200*3= 153600 -25600
2021 Purchases 230400*4.40= 1013760 1013760
COGAS 1192960 1167360
2021 Sales/COGS 51200*3.50 179200 230400*4.40= 1013760
179200*4.40= 788480 46080
Ending inventory 51200*4.40= 225280 51200*3= 153600 -71680
With Reference to above workings:
Journal entries to be passed for change from FIFO to LIFO
Tax Asset(25600*40%) 10240
Retained earnings(25600*60%) 15360
Inventory (decrease due to LIFO) 25600
(for 2020)
Tax Asset(46080*40%) 18432
Retained earnings(46080*60%) 27648
Inventory(decrease due to LIFO) 46080
(for 2021)
As the change is effective Jan 1, 2020, Income statement figures except net income & tax are for that year only & 2021 balance sheet figures are cumulative.
So, we need to change only the above 3 accounts in the balance sheet given.
As income tax has already been paid , we create a tax asset a/c like prepaid asset to be adjusted against future tax liability.
The re-casted Income statement &
Balance sheet are given as under:
Income for Years Ended
12/31/2020 12/31/2021
Sales 1152000 1728000
Less: COGS 672000 1013760
Other expenses 262400 389120
934400 1402880
Income before Income taxes 217600 325120
Income taxes (40%) 87040 130048
Net Income 130560 195072
Financial Position as at 12/31/2020 12/31/2021
Cash 166400 197120
Accounts receivable 128000 153600
Inventory 153600 153600
Other assets 217600 256000
Deferred tax asset 10240 28672
Total assets 675840 788992
Liabilities & Equity
Accounts payable 76800 102400
Other liabilities 102400 140800
Common stock 256000 256000
Retained earnings 240640 289792
Total liabilities and equity 675840 788992